OpenStudy (anonymous):

really need help A line has a rise of 6 and a slope of 1/20. What is the run?

OpenStudy (cwrw238):

1 / 20 = 6 / x find x

OpenStudy (anonymous):

or umm 120

OpenStudy (cwrw238):

no cross multiply x * 1 = 6* 20 x += ?

OpenStudy (anonymous):

whoa ok whoa i am so lost right now

OpenStudy (cwrw238):